# Optimal navigation in two-dimensional regular and turbulent flows

### v1

@@ V.4 Example 1: Steady vortex flow
First, we apply the RL algorithms to the navigation task in the steady vortex flow, see Section LABEL:sec:Ex_A. The flow domain is restricted to the square $0.3,1.2]^{2}$, which is discretized into $N_{s}=25^{2}$ uniform cells defining the states. The uncertainties in the initial and target points are controlled by $d_{A}=d_{B}=0.04$. Each episode has a maximum duration of $T_{max}=5$. The agents are trained for $n_{train}=10^{5}$ and then evaluated over $n_{eval}=5\cdot 10^{3}$ episodes. In Q-learning, we use a constant learning rate $\alpha=0.04$ and anneal the policy exploration parameter $\varepsilon$ from $1$ to $0.01$ by multiplying it by $0.9999$ after each episode. In the one-step actor-critic algorithm, the actor and critic learning rates are also constant and set to $\alpha_{p}=0.04$ and $\alpha_{w}=0.08$, respectively.
@@
The evolution of the total reward $r_{tot}=\sum_{t}r_{t}$ during training, averaged over a sliding window of $250$ episodes, is shown for both algorithms in Fig. LABEL:fig:10a. In each case, the total reward eventually stabilizes, indicating that the learning dynamics have reached a stationary regime and the algorithms have converged. Examples of trajectories obtained from the learned policies are presented in Fig. LABEL:fig:10b. For each method, we display ten realizations initialized uniformly at random within a disk of radius $d_{A}$ centered at $\bm{r}_{A}$. For reference, the optimal trajectory computed in Section LABEL:sec:OPC_A is shown alongside the RL-generated paths. The trajectories produced by Q-learning closely follow the optimal control solution; the small deviations that remain can be attributed primarily to the discretization inherent in our RL formulation. In contrast, the one-step actor-critic method yields trajectories that exhibit systematic deviations from the optimal path. These deviations are robust: they persist across independent training runs and remain even after varying the training hyperparameters. Apparently, this behavior can be related to the on-policy nature of one-step actor-critic algorithm, which may converge to local minimum of the policy-gradient landscape.
[…]
@@
The performance difference between the two methods is illustrated by the arrival-time statistics shown in Fig. LABEL:fig:10c. All evaluated trajectories reach the target, implying that both methods achieve a zero failure rate. However, the Q-learning agent achieves a slightly shorter mean travel time, $\langle T\rangle\approx 2.14$, compared with $\langle T\rangle\approx 2.18$ for the one-step actor-critic agent. Therefore, while both RL approaches successfully solve the navigation task, Q-learning provides a closer approximation to the optimal control solution and delivers better performance.
@@ V.5 Example 2: Steady sink and time-varying vortex flow
Next, we apply the Q-learning and one-step actor-critic algorithms to the navigation task in the steady sink and time-dependent vortex flow, see Section LABEL:sec:Ex_B. The training hyperparameters are identical to those used in the previous example, except that for the one-step actor-critic method we slightly increase the learning rates to $\alpha_{p}=0.05$ for the actor and $\alpha_{w}=0.1$ for the critic. The corresponding results are summarized in Fig. LABEL:fig:11.
Both algorithms converge successfully in this time-dependent external flow, as illustrated in Fig. LABEL:fig:11a, despite the fact that the state representation relies solely on the agent’s position. The trajectories produced by the learned policies are reasonably close to the optimal control solution derived in Section LABEL:sec:OPC_B, see Fig. LABEL:fig:11b. A visualization of the full time-dependent evolution is available in the videos provided in Ref. [repository]. All evaluated trajectories reach the target, demonstrating a zero failure rate for both approaches. The arrival-time statistics in Fig. LABEL:fig:11c indicate that the two algorithms perform comparably, with mean travel times of approximately $\langle T\rangle\approx 1.065$.
@@ V.6 Example 3: Taylor-Green vortices
The next example concerns the Taylor-Green flow described in Section LABEL:sec:Ex_C. The flow domain $0,2\pi]^{2}$ is discretized into $N_{s}=30^{2}$ uniform square cells that define the state space. The uncertainties in the initial and target positions are set to $d_{A}=d_{B}=0.2$. The maximum episode duration is given by $T_{\max}=2|\mathbf{r}_{B}-\mathbf{r}_{A}|/V_{0}$, where $V_{0}=0.1$ denotes the agent’s self-propulsion speed. The agents are trained for $n_{train}=2\cdot 10^{4}$ and then evaluated over $n_{eval}=5\cdot 10^{3}$ episodes. In Q-learning, we employ a constant learning rate of $\alpha=0.04$ and anneal the policy exploration parameter $\varepsilon$ from $1$ to $0.01$ by multiplying it by $0.9995$ after each episode. For the one-step actor-critic method, the actor and critic learning rates are also constant and set to $\alpha_{p}=0.005$ and $\alpha_{w}=0.01$, respectively.
@@
Figure LABEL:fig:12a shows the learning curves obtained for the Q-learning and one-step actor-critic algorithms. In both cases, the curves level off at relatively high total rewards, indicating that each method successfully learns an effective control policy. Representative trajectories generated by the learned policies are displayed in Fig. LABEL:fig:12b, alongside the optimal path computed in Section LABEL:sec:OPC_C. As illustrated, the Q-learning agent typically proceeds directly toward the target along a route close to the optimal one, although it may complete a single revolution around a nearby vortex when its initial state makes immediate escape impossible. The actor-critic agent performs less reliably: in addition to this initial revolution, it is sometimes captured by a second vortex, which forces an additional turn.
@@
These observations are further supported by the arrival-time statistics shown in Fig. LABEL:fig:12c. The first peak corresponds to near-optimal trajectories; the second arises from trajectories that include one extra loop within a Taylor-Green vortex; and the third reflects trajectories with two such additional rotations. Because of uncertainty in the initial position and the larger size of the target region, some trajectories terminate faster than the optimal-control prediction (LABEL:eq:opt-control-taylor-green). Note that all tested trajectories for both RL algorithms arrive at the target, yielding a zero failure rate. The appearance of the third peak leads to a slightly longer mean travel time for the actor-critic agent, $\langle T\rangle\approx 13.1$, compared with $\langle T\rangle\approx 12.9$ for the Q-learning agent. However, this difference is only a few percent, and overall both algorithms solve the task successfully.
@@
For comparison, in Ref. [repository] we have also implemented a trivial policy, where the steering angle selected by the agent at each $\Delta t$ is given by the action that points most directly toward the target among the $N_{a}=8$ different possibilities. In this case, the agent reaches the target only with a probability of $\approx 0.3$ (depending on its random starting position), and in the remaining attempts it ends up trapped in Taylor-Green vortices, from which this naive strategy does not allow to escape.
[…]
@@ V.7 Example 4: Snapshot of 2D Turbulence
[…]
sec:RL_2d_turb_snap
Next, we apply the RL algorithms to a navigation task in the steady velocity field obtained from a snapshot of two-dimensional turbulence, see Section LABEL:sec:Ex_D. The flow domain, $0,2\pi]^{2}$, is discretized into $N_{s}=30^{2}$ uniform square cells, which serve as the state space. Uncertainty in both the initial and target positions is characterized by $d_{A}=d_{B}=0.2$. The maximum episode duration is set to $T_{\max}=5|\mathbf{r}_{B}-\mathbf{r}_{A}|/V_{0}$, where $V_{0}$ is the agent’s self-propulsion speed. Agents are trained for $n_{train}=10^{5}$ episodes and evaluated over an additional $n_{eval}=5\times 10^{3}$ episodes. For Q-learning, we use a constant learning rate $\alpha=0.04$ and gradually reduce the policy exploration parameter $\varepsilon$ from $1$ to $0.01$ by multiplying it by $0.9999$ at the end of each episode. For the one-step actor-critic algorithm, the actor and critic learning rates are kept constant at $\alpha_{p}=0.005$ and $\alpha_{w}=0.01$, respectively.
Figure LABEL:fig:13a shows the evolution of the total reward throughout training. For both RL algorithms, the learning curves level off, indicating convergence. Representative trajectories generated by the learned policies are displayed in Fig. LABEL:fig:13b, together with the optimal control trajectory from Section LABEL:sec:OPC_D. As seen in the figure, the actor-critic agent’s trajectories closely track the optimal control path, whereas the Q-learning trajectories exhibit small deviations in the latter part of the journey. Consistently, the actor-critic agent achieves a shorter average travel time, $\langle T\rangle\approx 0.87$, compared with $\langle T\rangle\approx 0.96$ for the Q-learning agent, see Fig. LABEL:fig:13c.
This difference does not contradict the observation in Fig. LABEL:fig:13a, where the Q-learning agent attains a higher total reward at the end of training. The discrepancy arises because, during evaluation, the actor-critic agent selects the most probable actions from the learned policy, which yield higher rewards and consequently shorter travel times. Overall, all evaluated trajectories successfully reach the target, demonstrating a zero failure rate for both algorithms. Nonetheless, in the considered example the one-step actor-critic method provides a closer approximation to the optimal control solution and delivers better performance. Note also that the trivial policy implemented in Ref. [repository] completely fails to solve the considered navigation task, yielding a failure rate of $1$. Under this strategy, the agent eventually enters regions where the external flow cancels its self-propulsion, causing it to stall. As a result, it becomes effectively immobilized and fails to reach the target.
[…]
@@ V.8 Example 5: 2D Turbulence
[…]
sec:RL_2d_turb_time
In our final example, we consider navigation in a time-dependent 2d turbulent flow, see Section LABEL:sec:Ex_D. The Q-learning hyperparameters are identical to those used for the steady 2d turbulence snapshot, while for the one-step actor-critic algorithm we increase the learning rates to $\alpha_{p}=0.04$ for the actor and $\alpha_{w}=0.08$ for the critic. The obtained results are summarized in Fig. LABEL:fig:14.
Both methods converge, as shown by the learning curves in Fig. LABEL:fig:14a. The trajectories generated by their learned policies are similar to each other but diverge from the optimal-control solution presented in Section LABEL:sec:OPC_E, see Fig. LABEL:fig:14b. The discrepancies arise mainly in the second half of the path: the optimal control solution identifies a shorter and faster route through a region of weak external flow that the agent can overcome, whereas the RL policies exploit a strong background current pointing toward the goal. A corresponding video demonstration is provided in Ref. [repository]. Evidently, the shorter path requires more coordinated actions and is therefore more difficult for the RL agents to discover. The longer RL-derived path is more stable and still yields a high total reward, so both algorithms converge to it.
The performance difference between Q-learning and the actor-critic method is reflected in the arrival-time statistics in Fig. LABEL:fig:14c. All evaluated trajectories reach the target, indicating zero failure rate for both algorithms. However, the actor-critic agent exhibits a higher mean arrival time and a larger variance, implying that the Q-learning solution is slightly more efficient and robust. Overall, both methods successfully complete the navigation task in a complex 2d turbulent flow and demonstrate comparable performance.
Finally, Ref. [repository] also includes a trivial policy for this scenario. It fails to reach the target in most runs, with a failure rate close to $0.95$. Even when it succeeds, the travel time is very long and the resulting trajectories are not optimal.
@@ VI.1 Methodology
In two-dimensional turbulent flows, most of the kinetic energy is concentrated at large spatial scales, see Fig. LABEL:fig:1. In the inverse energy cascade range, $k<k_{f}$, where $k_{f}$ denotes the forcing wavenumber, the energy spectrum follows the scaling $E(k)\propto k^{-5/3}$. Consequently, the characteristic velocity associated with motions of size $r$ can be estimated as $v_{r}\sim(\epsilon r)^{1/3}$, where $\epsilon$ is the energy flux in the inverse cascade or, equivalently, the power injected into the system [boffetta2012two]. The characteristic timescale for velocity variations at scale $r$ then becomes $t_{r}\sim r/v_{r}\sim\epsilon^{-1/3}r^{2/3}$. Thus, the energetically dominant large scales evolve slowly and are relatively easy to predict, whereas small-scale fluctuations change rapidly and are inherently difficult to forecast with precision.
@@
We then train the agent using RL algorithms on this coarse-grained flow field and determine the resulting navigation strategies. Finally, we evaluate these strategies in the full velocity field to assess whether the agent can still successfully reach its target when exposed to the previously unseen small-scale fluctuations. Below, we examine two representative navigation scenarios: one in a steady snapshot of the 2d turbulent flow, and another in the fully time-dependent 2d turbulent field. In our examples, the forcing scale is fixed at $k_{f}=5.5$, and we progressively reduce the filtering scale, considering $k_{max}=\{10,5,3,2\}$.
[…]

### v2

@@ V.4 Example 1: Steady vortex flow
First, we apply the RL algorithms to the navigation task in the steady vortex flow, see Section III.1. The flow domain is restricted to the square $0.3,1.2]^{2}$, which is discretized into $N_{s}=25^{2}$ uniform cells defining the states. The uncertainties in the initial and target points are controlled by $d_{A}=d_{B}=0.04$. Each episode has a maximum duration of $T_{max}=5$. The agents are trained for $n_{train}=10^{5}$ and then evaluated over $n_{eval}=5\cdot 10^{3}$ episodes. In Q-learning, we use a constant learning rate $\alpha=0.04$ and anneal the policy exploration parameter $\varepsilon$ from $1$ to $0.01$ by multiplying it by $0.9999$ after each episode until it reaches the lower bound. In the one-step actor-critic algorithm, the actor and critic learning rates are also constant and set to $\alpha_{p}=0.04$ and $\alpha_{w}=0.08$, respectively.
@@
The evolution of the total reward $r_{tot}=\sum_{t}r_{t}$ during training, averaged over a sliding window of $250$ episodes, is shown for both algorithms in Fig. 10a. In each case, the total reward eventually stabilizes, indicating that the learning dynamics have reached a stationary regime and the algorithms have converged. Examples of trajectories obtained from the learned policies are presented in Fig. 10b. For each method, we display ten realizations initialized uniformly at random within a disk of radius $d_{A}$ centered at $\bm{r}_{A}$. For reference, the optimal trajectory computed in Section IV.2 is shown alongside the RL-generated paths. The trajectories produced by Q-learning closely follow the optimal control solution; the small deviations that remain can be attributed primarily to the discretization inherent in our RL formulation. In contrast, the one-step actor-critic method yields trajectories that exhibit systematic deviations from the optimal path. These deviations are robust: they persist across independent training runs and remain even after varying the training hyperparameters. Apparently, this behavior can be related to the on-policy nature of the one-step actor-critic algorithm, which may converge to local minimum of the policy-gradient landscape.
Figure 12: Example of Taylor-Green flow, see Section III.3. (a) Total reward, averaged over a sliding window of $250$ episodes, plotted against the number of training episodes. (b) Trajectories generated by the Q-learning (red) and one-step actor-critic (blue) agents, together with those obtained using the trivial policy (cyan) are shown alongside the optimal control trajectory (white). (c) Arrival-time histograms for the trained agents.
@@
The performance difference between the two methods is illustrated by the arrival-time statistics shown in Fig. 10c. All evaluated trajectories reach the target, implying that both methods achieve a zero failure rate. However, the Q-learning agent achieves a slightly shorter mean travel time, $\langle T\rangle\approx 2.14$, compared with $\langle T\rangle\approx 2.18$ for the one-step actor-critic agent. Therefore, while both RL approaches successfully solve the navigation task, Q-learning provides a closer approximation to the optimal control solution and delivers better performance, but the difference in mean travel time amounts to a few percent.
@@ V.5 Example 2: Steady sink and time-varying vortex flow
Next, we apply the Q-learning and one-step actor-critic algorithms to the navigation task in the steady sink and time-dependent vortex flow, see Section III.2. The training hyperparameters are identical to those used in the previous example, except that for the one-step actor-critic method we slightly increase the learning rates to $\alpha_{p}=0.05$ for the actor and $\alpha_{w}=0.1$ for the critic. The corresponding results are summarized in Fig. 11.
Both algorithms converge successfully in this time-dependent external flow, as illustrated in Fig. 11a, despite the fact that the state representation relies solely on the agent’s position. The trajectories produced by the learned policies are reasonably close to the optimal control solution derived in Section IV.3, see Fig. 11b. A visualization of the full time-dependent evolution is available in the videos provided in Ref. [repository]. All evaluated trajectories reach the target, demonstrating a zero failure rate for both approaches. The arrival-time statistics in Fig. 11c indicate that the two algorithms perform comparably, with mean travel times of approximately $\langle T\rangle\approx 1.065$.
@@ V.6 Example 3: Taylor-Green vortices
The next example concerns the Taylor-Green flow described in Section III.3. The flow domain $0,2\pi]^{2}$ is discretized into $N_{s}=30^{2}$ uniform square cells that define the state space. The uncertainties in the initial and target positions are set to $d_{A}=d_{B}=0.2$. The maximum episode duration is given by $T_{\max}=2|\mathbf{r}_{B}-\mathbf{r}_{A}|/V_{0}$, where $V_{0}=0.1$ denotes the agent’s self-propulsion speed. The agents are trained for $n_{train}=2\cdot 10^{4}$ and then evaluated over $n_{eval}=5\cdot 10^{3}$ episodes. In Q-learning, we employ a constant learning rate of $\alpha=0.04$ and anneal the policy exploration parameter $\varepsilon$ from $1$ to $0.01$ by multiplying it by $0.9995$ after each episode until the lower bound is reached. For the one-step actor-critic method, the actor and critic learning rates are also constant and set to $\alpha_{p}=0.005$ and $\alpha_{w}=0.01$, respectively.
@@
Figure 12a shows the learning curves obtained for the Q-learning and one-step actor-critic algorithms. In both cases, the curves level off at relatively high total rewards, indicating that each method successfully learns an effective control policy. Representative trajectories generated by the learned policies are displayed in Fig. 12b, alongside the optimal control path computed in Section IV.4. As illustrated, the Q-learning agent typically proceeds directly toward the target along a route close to the optimal one, although it may complete a single revolution around a nearby vortex when its initial state makes immediate escape impossible. The actor-critic agent performs less reliably: in addition to this initial revolution, it is sometimes captured by a second vortex, which forces an additional turn.
@@
These observations are further supported by the arrival-time statistics shown in Fig. 12c. The first peak corresponds to near-optimal trajectories; the second arises from trajectories that include one extra loop within a Taylor-Green vortex; and the third reflects trajectories with two such additional rotations. Because of uncertainty in the initial position and the larger size of the target region, some trajectories terminate faster than the optimal-control prediction (25). Note that all tested trajectories for both RL algorithms arrive at the target, yielding a zero failure rate. The appearance of the third peak leads to a slightly longer mean travel time for the actor-critic agent, $\langle T\rangle\approx 13.1$, compared with $\langle T\rangle\approx 12.9$ for the Q-learning agent. However, this difference is only a few percent, and overall both algorithms solve the task successfully.
@@
For comparison, we have also implemented a trivial policy, where the steering angle selected by the agent at each $\Delta t$ is given by the action that points most directly toward the target among the $N_{a}=8$ different possibilities. In this case, the agent reaches the target with a probability of only $\approx 0.3$, depending on its random starting position. In the remaining trials, the agent becomes trapped in Taylor–Green vortices, from which this naive strategy cannot escape, as illustrated by the cyan trajectories in Fig. 12b.
Figure 13: Example of 2d turbulence snapshot, see Section III.4. (a) Total reward, averaged over a sliding window of $250$ episodes, plotted against the number of training episodes. (b) Trajectories generated by the Q-learning (red) and one-step actor-critic (blue) agents, together with those obtained using the trivial policy (cyan) are shown alongside the optimal control trajectory (white). (c) Arrival-time histograms for the trained agents.
@@ V.7 Example 4: Snapshot of 2D Turbulence
Next, we apply the RL algorithms to a navigation task in the steady velocity field obtained from a snapshot of two-dimensional turbulence, see Section III.4. The flow domain $0,2\pi]^{2}$ is discretized into $N_{s}=30^{2}$ uniform square cells, which serve as the state space. Uncertainty in both the initial and target positions is characterized by $d_{A}=d_{B}=0.2$. The maximum episode duration is set to $T_{\max}=5|\mathbf{r}_{B}-\mathbf{r}_{A}|/V_{0}$, where $V_{0}$ is the agent’s self-propulsion speed. Agents are trained for $n_{train}=10^{5}$ episodes and evaluated over an additional $n_{eval}=5\cdot 10^{3}$ episodes. For Q-learning, we use a constant learning rate $\alpha=0.04$ and gradually reduce the policy exploration parameter $\varepsilon$ from $1$ to $0.01$ by multiplying it by $0.9999$ at the end of each episode until the lower bound is reached. For the one-step actor-critic algorithm, the actor and critic learning rates are kept constant at $\alpha_{p}=0.005$ and $\alpha_{w}=0.01$, respectively.
Figure 13a shows the evolution of the total reward throughout training. For both RL algorithms, the learning curves level off, indicating convergence. Representative trajectories generated by the learned policies are displayed in Fig. 13b, together with the optimal control trajectory from Section IV.5. As seen in the figure, the actor-critic agent’s trajectories closely track the optimal control path, whereas the Q-learning trajectories exhibit small deviations in the latter part of the journey. Consistently, the actor-critic agent achieves a shorter average travel time, $\langle T\rangle\approx 0.87$, compared with $\langle T\rangle\approx 0.96$ for the Q-learning agent, see Fig. 13c.
This difference does not contradict the observation in Fig. 13a, where the Q-learning agent attains a higher total reward at the end of training. The discrepancy arises because, during evaluation, the actor-critic agent selects the most probable actions from the learned policy, which yield higher rewards and consequently shorter travel times. Overall, all evaluated trajectories successfully reach the target, demonstrating a zero failure rate for both algorithms. Nonetheless, in the considered example the one-step actor-critic method provides a closer approximation to the optimal control solution and delivers better performance. Note also that the trivial policy completely fails to solve the considered navigation task, yielding a failure rate of unity. Under this strategy, the agent eventually enters regions where the external flow cancels its self-propulsion, causing it to stall, see the cyan trajectories in Fig. 13b. As a result, it becomes effectively immobilized and fails to reach the target.
Figure 14: Example of time-dependent 2d turbulent flow, see Section III.4. (a) Total reward, averaged over a sliding window of $250$ episodes, plotted against the number of training episodes. (b) Trajectories generated by the Q-learning (red) and one-step actor-critic (blue) agents, together with those obtained using the trivial policy (cyan) are shown alongside the optimal control trajectory (white). (c) Arrival-time histograms for the trained agents.
@@ V.8 Example 5: 2D Turbulence
In our final example, we consider navigation in a time-dependent 2d turbulent flow, see Section III.4. The training hyperparameters are identical to those used for the steady 2d turbulence snapshot, except the tuned learning rates. Specifically, now the Q-learning algorithm uses a learning rate of $\alpha=0.01$, while the one-step actor-critic method employs learning rates of $\alpha_{p}=0.04$ for the actor and $\alpha_{w}=0.08$ for the critic. The obtained results are summarized in Fig. 14.
Both methods converge, as indicated by the learning curves in Fig. 14a. However, the learned policies differ substantially, which is reflected in the distinct trajectories followed by the trained agents, see Fig. 14b. The Q-learning agent follows a path that is close to the solution obtained within optimal control theory in Section IV.6. The latter part of this trajectory passes through a region of weak background flow, which the agent is able to overcome by executing well-coordinated maneuvers. As can be observed in the demonstration video provided in Ref. [repository], the Q-learning agent selects suboptimal motion directions near the very end of the episode, resulting in a noticeable increase in the mean travel time $\langle T\rangle\approx 0.096$ compared to $\langle T\rangle\approx 0.07$ for the optimal control solution, see Eq. (30).
The actor-critic agent, by contrast, follows a different route. It identifies a strong background current directed toward the target and successfully exploits it. This strategy yields a mean travel time of $\langle T\rangle\approx 0.097$, which is comparable to the performance of the Q-learning agent. Both methods exhibit a zero failure rate, and the corresponding arrival-time statistics are presented in Fig. 14c. Qualitatively, the observed behavior of the agents is similar to the navigation task in a steady vortex flow, where different RL methods also discover distinct trajectories with comparable performance, see Section V.4. In both examples, the trajectories discovered by the Q-learning agent more closely resemble the theoretic optimal control solution to the navigation problem.
Finally, we examine a trivial policy for the 2d turbulent flow example. This policy fails to reach the target in the majority of runs, with a failure rate of approximately $0.95$. Even in the rare cases where it succeeds, the travel time is excessively long and the resulting trajectories are far from optimal. Figure 14b shows representative trajectories for this policy, depicted in cyan. The low self-propulsion speed of the agent prevents it from overcoming the external currents and reaching the target.
@@ VI.1 Methodology
In two-dimensional turbulent flows, most of the kinetic energy is concentrated at large spatial scales, see Fig. 1. In the inverse energy cascade range, $k<k_{f}$, where $k_{f}$ denotes the forcing wavenumber, the energy spectrum follows the scaling $E(k)\propto k^{-5/3}$. Consequently, the characteristic velocity associated with motions of size $r$ can be estimated as $v_{r}\sim(\epsilon r)^{1/3}$, where $\epsilon$ is the energy flux in the inverse cascade or, equivalently, the power injected into the system [boffetta2012two]. The characteristic timescale for velocity variations at scale $r$ then becomes $t_{r}\sim r/v_{r}\sim\epsilon^{-1/3}r^{2/3}$. Thus, the energetically dominant large scales evolve slowly and are relatively easy to predict, whereas small-scale fluctuations change rapidly and are inherently difficult to forecast with precision.
@@
We then train the agent using RL algorithms on this coarse-grained flow field and determine the resulting navigation strategies. Finally, we evaluate these strategies in the full velocity field to assess whether the agent can still successfully reach its target when exposed to the previously unseen small-scale fluctuations. Below, we examine two representative navigation scenarios: one in a steady snapshot of the 2d turbulent flow, and another in the fully time-dependent 2d turbulent field. In our examples, the forcing scale is fixed at $k_{f}=5.5$, and we progressively reduce the filtering scale, considering $k_{max}=\{256,10,5,3,2\}$. Note that $k_{max}=256$ corresponds to the fully resolved DNS velocity field.
The performance of the agents is quantified using two primary metrics: the fraction of attempts that successfully reach the target and the mean travel time of those successful attempts. Both quantities are analyzed as functions of the normalized filtering scale $k_{\max}/k_{f}$. For each value of $k_{\max}$ and for each RL algorithm, we perform $n=10$ independent training runs and report averages over these runs to reduce variability due to stochastic training effects. The mean success probability is defined as
$\langle p\rangle=\frac{1}{n}\sum_{i=1}^{n}p_{i},$ (53)
where $p_{i}$ denotes the success probability obtained in the $i$-th training run. The mean travel time, computed over successful trajectories only, is given by
$\langle T\rangle=\sum_{i=1}^{n}g_{i}\langle T\rangle_{i},\qquad g_{i}=\frac{p_{i}}{\sum_{k=1}^{n}p_{k}},$ (54)
where $\langle T\rangle_{i}$ is the mean travel time of successful episodes in run $i$, and the weights $g_{i}$ ensure that runs with higher success rates contribute proportionally more to the overall average. In addition to the mean values, we report the estimated standard deviations of both metrics across independent training runs. These fluctuations provide a quantitative measure of the robustness and reproducibility of agent performance with respect to stochasticity in the training procedure.
Figure 15: Trajectories produced by the Q-learning (red) and one-step actor-critic (blue) agents are shown together with the optimal control trajectory (white) for the 2d turbulence snapshot. The background depicts the coarse-grained velocity field, retaining only modes with wavenumbers $k\leq k_{max}$, which was used during training. All trajectories, however, were evaluated in the full (unfiltered) velocity field.